### v1

@@ Abstract
Identifying (a) systemic barriers to quality healthcare access and (b) key indicators of care efficacy in the United States remains a significant challenge. To improve our understanding of regional disparities in care delivery, we introduce a novel application of curvature, a geometrical-topological property of networks, to Physician Referral Networks. Our initial findings reveal that Forman-Ricci and Ollivier-Ricci curvature measures, which are known for their expressive power in characterizing network structure, offer promising indicators for detecting variations in healthcare efficacy while capturing a range of significant regional demographic features. We also present apparent [^1] Access our interactive Datasette prototype at https://apparent.topology.rocks/.  [^2] Access our associated analysis repository at https://github.com/aidos-lab/apparent., an open-source tool that leverages Ricci curvature and other network features to examine correlations between regional Physician Referral Networks structure, local census data, healthcare effectiveness, and patient outcomes.
@@ 1 Introduction
In the rapidly evolving field of healthcare management, the analysis of medical claims data has become an essential component for improving the quality and equity of healthcare services. The nature of care delivery in the United states is heavily influenced by its fragmentation—care is often spread across multiple disconnected providers (e.g., primary-care physicians, specialists). Settings with greater care fragmentation have been shown to inhibit effective communication and coordination between care team members, thus contributing to higher costs and lower quality of treatment [13, 33, 21, 1, 7].
@@
Despite the well-understood impacts of fragmentation, there are still few quantitative tools that can capture the mechanisms of care delivery networks at scale [14]. Standard analyses of local infrastructure features, often executed using tabular data, are limited in their ability to distill complex dynamics between physicians. Graph representations and tools from network analysis have thus been adopted in the literature to provide more sophisticated models for care delivery [19, 29, 26, 2, 3, 4, 6, 12, 14, 22, 25, 28, 6], permitting an understanding of the data both on the local and the global level. Increased access to rich data has placed an emphasis on analyzing Physician Referral Networks, which encode patient sharing between physicians across different regions in the United States. Physician Referral Networks are typically modeled using administrative claims data (often from Medicare) to map patient-flow patterns, elucidating more complex dynamics that underpin local care delivery. Their study has proved fruitful for enhancing care coordination, reducing costs, improving patient outcomes, and understanding social determinants of health [10, 15, 17, 24].
@@
Our work aims to introduce novel foundational tools to the study of Physician Referral Networks. In particular, we focus on curvature-based measures for illuminating structural properties of a graph. Ricci curvature, a well-known concept in differential geometry, has recently gained traction for various applications in the field of machine learning [34, 8, 9]. In fact, positive Ricci curvature in a network was shown to correspond to regions of high connectivity–-–which, in the context of care delivery, could represent tighter-knit groups with the capacity for better coordination and communication. Negative curvature, by contrast, is known to capture information bottlenecks in message-passing neural networks [36], and has the potential to indicate less robust regions in a referral network, prone to information loss, miscommunication, and potentially lower standards of care.
@@
In addition to strong theoretical foundations and promising interpretations in the context of care delivery, Ricci curvature measures also stand out against standard network descriptors (e.g., assortativity, centrality, clustering coefficents) for their ability to precisely describe the structure of a network and its flexibility in easily switching between edge-level, node-level, or network-level aggregations to suit the analysis at hand [35]. Thus, we aim to bolster foundational analysis of Physician Referral Networks with the following key contributions:
@@ 2 Background
Research on Physician Referral Networks has evolved significantly, transitioning from small-scale empirical studies to large-scale analyses enabled by the advent of large-scale data access [30, 31, 20, 24, 5, 14]. This progression has expanded the scope and depth of understanding regarding the structure and dynamics of care delivery networks. Early empirical studies, despite being constrained to individual care delivery networks and often involving only several thousand physicians, provided essential insights into the conceptualization of Physician Referral Networks and their ability to characterize healthcare delivery [10, 24]. Recently, the growing availability of large-scale administrative data, particularly from electronic medical records and insurance claims has allowed for more robust studies on Physician Referral Networks across the country. This work examines quantitative network properties to glean insights on care coordination, utilization, and cost as well as relationships between network structure and economic/clinical outcomes [15, 19, 29, 12].
However, there remain significant challenges to effectively and efficiently extracting relevant features from these complex data structures. Key obstacles include the high dimensionality of network data, the computational complexity of analyzing large-scale networks, and the difficulty in interpreting network measures in clinical contexts. This creates a crucial need for new analysis approaches. Bridging the gap between foundational theory and contemporary measurement has already shown promising results for care delivery analysis [14, 22, 28]. By introducing Ricci curvature, we aim to further expand this frontier, leveraging a rich mathematical tradition in differential geometry to enhance our collective interpretations when evaluating these intricate healthcare systems.
@@
Table 1: Data Release Summary. This table summarizes the SQL database tables contained in apparent, along with a description of their different features their importance for characterizing local Physician Referral Networks.
@@ 3 Data Description
We analyze Medicare claims data in the United States from 2014–2017, consisting of 66,3541 total physicians, with approximately 10,000,000 patient-sharing records per year across 3404 Hospital Service Areas (HSAs). These data were made publicly available by CareSet, [^3] https://www.careset.com/ a health care data consultancy, in collaboration with the Center for Medicare and Medicaid Services (CMS), the United States government agency that runs Medicare, and have been widely used in prior research on physician networks [16, 11, 18, 23]. We supplemented these data with additional information from several sources. Moreover, we obtained basic data on the physicians included in our networks (e.g., practice locations, specialty types) from the National Plan and Provider Enumeration System (NPPES). We also collected data on local health care quality and spending from the Dartmouth Institute for Health Policy and Clinical Practice.
@@
Health care delivery in the United States tends to be highly geographically localized, and there is known to be substantial variation in both health care practices and outcomes across regions. To capture this variation, we map physician networks separately by region, focusing on Hospital Service Areas (HSAs; as defined by the Dartmouth Atlas). HSAs have been widely used in prior research on health care delivery, and are designed to correspond roughly to local markets for hospital services. We position physicians within HSAs by geocoding their practice locations; as such, our network maps consist of shared patients among the physicians practicing within a particular HSA. Finally, we map networks separately by year, for each of the four years in our study window, out of recognition that the structure of Physician Referral Networks may change over time.
@@
To protect patient privacy and comply with CMS regulations, we only include physician pairs in our data if they shared at least 11 patients within a specific time period. While this thresholding approach necessarily results in some data loss, previous research suggests that connections between physicians based on few shared patients are often not considered significant by the physicians themselves [5]. The most meaningful professional relationships typically involve a higher number of shared patients and align more closely with survey data. Therefore, this required minimum threshold is actually beneficial, as it likely produces networks that more accurately reflect real-world professional relationships.
@@ 4.1 Building Physician Referral Networks
Our goal is to capture patient flow within an HSA during the course of a year. Each provider in the HSA becomes a node in the graph, and edges are are assigned between physicians when sufficiently many patients are shared via referrals. We provide some real examples in Figure 1 that arise from HSAs in different regions of the country (Southern California vs. Atlanta). Our graph representations are built directly from the local_physician_interactions table described in Table 1.
@@
Figure 1: Visualizing Physician Referral Networks. Above, we visualize networks from regions with known differences in care delivery. Figure 1(a) shows a referral network from Santa Ana, California, exhibiting a high quality of care, while Figure 1(b) comes from Snellville, Georgia. As made available in apparent, we combine traditional network features with Ricci Curvature to elucidate the structural properties of care delivery in each region: Nodes are sized and colored by centrality and degree while edges are sized and colored by number of referrals and $\kappa_{\text{OR}}$, respectively. Interpretation: Larger yellow nodes are physicians that play a prominent role in regional care delivery by dictating patient flow. Blue-gray edges (negative curvature) are potential bottlenecks for patients finding the best provider, while orange edges (positive curvature) indicate strong communication between physicians and well-established patient flow.
@@ 4.2.1 Forman-Ricci Curvature
Here, $d_{i}$ and $d_{j}$ denote the degrees of nodes $i$ and $j$, respectively, and |#$|\#_{\Delta}|$ represents the number of triangles (i.e., 3-cliques) that include the edge $(i,j)$. In Physician Referral Networks, $d_{i}$ and $d_{j}$ correspond to the number of direct connections (referrals) each physician has, while |#$|\#_{\Delta}|$ indicates the number of mutual connections between the two physicians. A higher FRC value suggests a stronger collaborative relationship, characterized by more mutual connections and shared professional interactions. In general FRC values are unbounded—in large networks as (see Figure 4), edges can take large positive or large negative values. As an intuitive example for an edge with very negative $\kappa_{\text{FR}}$, imagine two communities modeled as fully connected graphs that are connected by a single edge. Each community can be made to have arbitrarily large degrees ($d_{i},d_{j}$ in Equation 1), but the edge between them cannot have mutual connections (i.e., |#$|\#_{\Delta}|=0$).
@@ 4.2.2 Ollivier-Ricci Curvature
In the context of physician networks, $\mu_{i}$ and $\mu_{j}$ can be understood as distributions of patient referrals or collaborations within the neighborhoods of physicians $i$ and $j$. A higher ORC value indicates that the neighborhoods of the two nodes are more similar, suggesting more cohesive collaboration patterns. For extreme examples or positive and negative ORC, see Figure 2. Despite its enhanced expressiveness in capturing distinct structural properties of networks, even for challenging tasks such as distinguishing Strongly-Regular and Rook-Shrikande graphs [35], ORC is computationally expensive for very large graphs. In contrast, FRC offers a more scalable alternative. If FRC can capture meaningful signals, it provides a practical advantage as it can be computed efficiently, even for the largest Physician Referral Networks in the United States.
@@ 5 Results
Our interactive database contains information on approximately 13,500 physician networks (bounded by HSAs), sourced over the years 2014–2017. Along with approximately 20 network descriptors, including statistics for network assortativity, clustering coefficients, density centrality, Forman-Ricci Curvature, and Ollivier-Ricci Curvature, we also provide infrastructure for querying rich metadata for these networks. Table 1 provides a summary of the available tables and their basic descriptions.
@@
Figure 3: Regional Distribution of Physician Referral Networks. Here we visualize the distribution of Physician Referral Networks (one per Health Service Area, HSA) and physician counts (nodes) by state, color-coded by region. This figure highlights the relative size and density of Physician Referral Networks across the United States in 2017.
Additionally, we provide distributions of Forman-Ricci and Ollivier-Ricci curvatures values for network edges, grouped by state and region in Figures 4 and 5. As we observe from the distribution, ORC values are constrained to a much tighter range, with the most negatively curved edges in our distribution approaching a value of $-2$, while all edge curvatures are bounded above by $+1$. In comparison, Forman-Ricci, due to its dependence on node degree and the number of triangles, can produce very large curvature values, both negative and positive. Given the known trade-off between locality, expressivity, and efficiency between Forman-Ricci and Ollivier-Ricci curvature [34], further analysis is required to understand which descriptors are most useful for specific healthcare management tasks. Despite clear agreement between various classical network science features (e.g, Figure 6), Ricci curvature measures are known to be much more expressive in capturing network structure. Moreover, there is precedence for alleviating bottlenecks in graph neural networks by rewiring negatively curved edges in a network [36]. This provides a novel way to potentially “rewire” care delivery networks using these measures in order to improve healthcare quality in regions that suffer from care fragmentation.
Figure 4: Regional Distribution of Forman-Ricci Curvature. This figure displays a wide range of inter- and intra-regional differences in care delivery structure as measured by $\kappa_{\text{FR}}$. We plot curvature values for each edge (provider connection) for all 2017 Physician Referral Networks, separated by state and colored by region.
Figure 5: Regional Distribution of Ollivier-Ricci Curvature This figure displays the regional differences in care delivery structure as measured by $\kappa_{\text{OR}}$. We plot curvature values for each edge (provider connection) for all 2017 Physician Referral Networks, separated by state and colored by region. In comparison with $\kappa_{\text{FR}}$, (see Figure 4), we see much smaller variance in $\kappa_{\text{OR}}$ values which is expected based on their construction in Equation 1 and Equation 2. Although $\kappa_{\text{OR}}$ is known to be a more expressive measure for capturing fine-grained network structure, although it can be much more expensive to compute for large networks.
@@
Figure 6: Comparing Standard Network Features and Ricci Curvature. For a select set of states in the Philidelphia (green), Chicago (purple), and Denver (gray) HSA regions, we visualize distributions of clustering coefficents (CC) in comparison to $\kappa_{\text{OR}}$. These are different measures (with clearly different distributions) for capturing connectivity in a network. In the context of care delivery positive CC and $\kappa_{\text{OR}}$ values together could indicate high functioning sub regions of a Physician Referral Network.
As part of our preliminary analysis, we compared Forman-Ricci curvature with various population statistics, including Medicare enrollment, total mortality, and the proportion of non-white populations. Consistent with Figure 7, our findings reveal that large networks exhibiting low average curvature—indicating many potential bottlenecks in care delivery—correlate with high Medicare enrollment, elevated mortality rates, and higher proportions of Black and Hispanic populations. These results (which can be reproduced using our novel apparent interactive tool) suggest that negatively curved edges are indicative of structural properties in physician networks commonly found in less affluent HSAs.

### v2

@@ Abstract
Identifying (a) systemic barriers to quality healthcare access and (b) key indicators of care efficacy in the United States remains a significant challenge. To improve our understanding of regional disparities in care delivery, we introduce a novel application of curvature, a geometrical-topological property of networks, to Physician Referral Networks. Our initial findings reveal that Forman-Ricci and Ollivier-Ricci curvature measures, which are known for their expressive power in characterizing network structure, offer promising indicators for detecting variations in healthcare efficacy while capturing a range of significant regional demographic features. We also present apparent, an open-source tool that leverages Ricci curvature and other network features to examine correlations between regional Physician Referral Networks structure, local census data, healthcare effectiveness, and patient outcomes.
@@ 1 Introduction
In the rapidly evolving field of healthcare management, the analysis of medical claims data has become an essential component for improving the quality and equity of healthcare services. The nature of care delivery in the United States is heavily influenced by its fragmentation—care is often spread across multiple disconnected providers (e.g., primary-care physicians, specialists). Settings with greater care fragmentation have been shown to inhibit effective communication and coordination between care team members, thus contributing to higher costs and lower quality of care [14, 34, 22, 1, 7].
@@
Despite the well-understood impacts of fragmentation, there are still few quantitative tools that can capture the mechanisms of care delivery networks at scale [15]. Standard analyses of local infrastructure features, often executed using tabular data, are limited in their ability to distill complex dynamics among physicians. Graph representations and tools from network analysis have thus been adopted in the literature to provide more sophisticated models for care delivery [20, 30, 27, 2, 3, 4, 6, 13, 15, 23, 26, 29, 6], permitting an understanding of the data both on the local and the global level. Increased access to medical claims data has placed an emphasis on analyzing Physician Referral Networks, which encode patient sharing between physicians across different regions in the United States. Physician Referral Networks are typically modeled using administrative claims data (often from Medicare) to map patient-flow patterns, elucidating more complex dynamics that underpin local care delivery. Their study has proved fruitful for enhancing care coordination, reducing costs, improving patient outcomes, and understanding social determinants of health [11, 16, 18, 25].
@@
Our work aims to introduce novel foundational tools to the study of Physician Referral Networks. In particular, we focus on curvature-based measures for capturing structural properties of a graph. Ricci curvature, a well-known concept in differential geometry, has recently gained traction for various applications in the field of machine learning [35, 9, 10]. In fact, positive Ricci curvature in a network was shown to correspond to regions of high connectivity–-–which, in the context of care delivery, could represent tighter-knit groups with the capacity for better coordination and communication. Negative curvature, by contrast, is known to capture information bottlenecks in message-passing neural networks [37], and has the potential to indicate less robust regions in a referral network that are prone to information loss, miscommunication, and potentially lower standards of care.
@@
In addition to strong theoretical foundations and promising interpretations in the context of care delivery, Ricci curvature measures also stand out against standard network descriptors (e.g., assortativity, centrality, clustering coefficients) due to their ability to precisely describe the structure of a network and their flexibility in easily switching between edge-level, node-level, or network-level aggregations to suit the analysis at hand [36]. Thus, we aim to bolster foundational analysis of Physician Referral Networks with the following key contributions:
@@ 2 Background
Research on Physician Referral Networks has evolved significantly, transitioning from small-scale empirical studies to large-scale analyses enabled by the advent of large-scale data access [31, 32, 21, 25, 5, 15]. This progression has expanded the scope and depth of understanding regarding the structure and dynamics of care delivery networks. Early empirical studies, despite being constrained to individual care delivery networks and often involving only several thousand physicians, provided essential insights into the conceptualization of Physician Referral Networks and their ability to characterize healthcare delivery [11, 25]. Recently, the growing availability of large-scale administrative data, particularly from electronic medical records and insurance claims, has allowed for more robust studies on Physician Referral Networks across the country. This work examines quantitative network properties to glean insights on care coordination, utilization, and cost, as well as relationships between network structure and economic/clinical outcomes [16, 20, 30, 13].
However, significant challenges remain in effectively and efficiently extracting relevant features from these complex data structures. Key obstacles include the high dimensionality of network data, the computational complexity of analyzing large-scale networks, and the difficulty in interpreting network measures in clinical contexts. This creates a crucial need for new analytical approaches. Bridging the gap between foundational theory and contemporary measurement has already shown promising results for care delivery analysis [15, 23, 29]. By introducing Ricci curvature, we aim to further expand this frontier, leveraging a rich mathematical tradition in differential geometry to enhance our collective interpretations when evaluating these intricate healthcare systems.
@@
Table 1: Data Release Summary. We provide an overview of the tables that comprise our original database, along with high level descriptions of their features and relevance for characterizing local Physician Referral Networks. Up-to-date table structure, data summaries, and feature descriptions will be maintained within our prototype, apparent.
@@ 3 Data Description
We analyze Medicare claims data in the United States from 2014–2017, consisting of 663,541 total physicians, with approximately 10,000,000 patient-sharing records per year across 3404 Hospital Service Areas (HSAs). These data were made publicly available by CareSet a healthcare data consultancy, in collaboration with the Center for Medicare and Medicaid Services (CMS), the U.S. government agency that administers Medicare, and have been widely used in prior research on physician networks [17, 12, 19, 24]. As outlined in Table 1, we supplement physician interactions with data from several sources including basic data on the physicians included in our networks (e.g., practice locations, specialty types) from the National Plan and Provider Enumeration System (NPPES) as well as metrics on local healthcare quality and spending from the Dartmouth Institute for Health Policy and Clinical Practice.
@@
Healthcare delivery in the United States tends to be highly localized, with substantial variation in both healthcare practices and outcomes across regions. To capture this variation, we map physician networks separately by region, focusing on Hospital Service Areas (HSAs; as defined by the Dartmouth Atlas). HSAs have been widely used in prior research on healthcare delivery and are designed to correspond roughly to local markets for hospital services. We position physicians within HSAs by geocoding their practice locations; as such, our network maps consist of shared patients among the physicians practicing within a particular HSA. Finally, we study networks separately by year, for each of the four years in our study window, recognizing that the structure of Physician Referral Networks may change over time.
@@
To protect patient privacy and comply with CMS regulations, we only include physician pairs in our data if they shared at least 11 patients within a specific time period [8]. While this thresholding approach necessarily results in some data loss, previous research suggests that connections between physicians based on few shared patients are often not considered significant by the physicians themselves [5]. The most meaningful professional relationships typically involve a higher number of shared patients and align more closely with survey data. Therefore, this required minimum threshold is actually beneficial, as it likely produces networks that more accurately reflect real-world professional relationships.
@@ 4.1 Building Physician Referral Networks
Our goal is to capture patient flow within an HSA during the course of a year. Each provider in the HSA becomes a node in the graph, and edges are assigned between physicians when a sufficient number of patients are shared via referrals. We provide some real examples in Figure 1 that arise from HSAs in different regions of the country (Southern California vs. Atlanta). Our graph representations are built directly from the local_physician_interactions table (see Table 1).
@@
Figure 1: Visualizing Physician Referral Networks. Above, we visualize networks from regions with known differences in care delivery. Figure 1(a) shows a referral network from Santa Ana, California, exhibiting a higher quality of care, while Figure 1(b) comes from Snellville, Georgia. As made available in apparent, we combine traditional network features with Ricci Curvature to elucidate the structural properties of care delivery in each region: Nodes are sized and colored by centrality and degree while edges are sized and colored by number of referrals and $\kappa_{\text{OR}}$, respectively. Interpretation: Larger yellow nodes are physicians that play a prominent role in regional care delivery by dictating patient flow. Blue-gray edges (negative curvature) are potential bottlenecks for patients finding the best provider, while orange edges (positive curvature) indicate strong communication between physicians and well-established patient flow.
@@ 4.2.1 Forman-Ricci Curvature
Here, $d_{i}$ and $d_{j}$ denote the degrees of nodes $i$ and $j$, respectively, and |#$|\#_{\Delta}|$ represents the number of triangles (i.e., 3-cliques) that include the edge $(i,j)$. In Physician Referral Networks, $d_{i}$ and $d_{j}$ correspond to the number of direct connections (referrals) each physician has, while |#$|\#_{\Delta}|$ indicates the number of mutual connections between the two physicians. A higher FRC value suggests a stronger collaborative relationship, characterized by more mutual connections and shared professional interactions. In general FRC values are unbounded—in large networks edges can take large positive or large negative values (see Figure 4). As an intuitive example for an edge with very negative $\kappa_{\text{FR}}$, imagine two communities modeled as fully connected graphs that are connected by a single edge. Each community can be made to have arbitrarily large degrees ($d_{i},d_{j}$ in Equation 1), but the edge between them cannot have mutual connections (i.e., |#$|\#_{\Delta}|=0$).
Figure 3: Regional Distribution of Physician Referral Networks. Here we visualize the distribution of Physician Referral Networks in 2017 (one per Health Service Area, HSA) alongside physician counts (nodes) by state, color-coded by region. This figure highlights the relative size and density of Physician Referral Networks and establishes a baseline for structural differences in care coordination across the United States.
@@ 4.2.2 Ollivier-Ricci Curvature
In the context of physician networks, $\mu_{i}$ and $\mu_{j}$ can be understood as distributions of patient referrals or collaborations within the neighborhoods of physicians $i$ and $j$. A higher ORC value indicates that the neighborhoods of the two nodes are more similar, suggesting more cohesive collaboration patterns. For extreme examples or positive and negative ORC, see Figure 2. Despite its enhanced expressiveness in capturing distinct structural properties of networks, even for challenging tasks such as distinguishing Strongly-Regular and Rook-Shrikhande graphs [36], ORC is computationally expensive for very large graphs. In contrast, FRC offers a more scalable alternative. If FRC can capture meaningful signals, it provides a practical advantage as it can be computed efficiently, even for the largest Physician Referral Networks in the United States.
@@ 5 Results
Our interactive database contains information on approximately 13,500 physician networks (bounded by HSAs), sourced over the years 2014–2017. Along with approximately 20 network descriptors, including statistics for network assortativity, clustering coefficients, density, centrality, Forman-Ricci Curvature, and Ollivier-Ricci Curvature, we also provide infrastructure for querying rich metadata for these networks. Table 1 provides a summary of the available tables and their basic descriptions.
@@
Figure 4: Regional Distribution of Forman-Ricci Curvature. This figure displays a wide range of inter- and intra-regional differences in care delivery structure as measured by $\kappa_{\text{FR}}$. We plot curvature values for each edge (provider connection) for all 2017 Physician Referral Networks.
Figure 5: Regional Distribution of Ollivier-Ricci Curvature This figure displays the regional differences in care delivery structure as measured by $\kappa_{\text{OR}}$. We plot curvature values for each edge (provider connection) for all 2017 Physician Referral Networks, separated by state and colored by region. In comparison with $\kappa_{\text{FR}}$ (see Figure 4), we see much smaller variance in $\kappa_{\text{OR}}$ values which is expected based on their construction in Equation 1 and Equation 2. Although $\kappa_{\text{OR}}$ is known to be a more expressive measure for capturing fine-grained network structure, although it can be much more expensive to compute for large networks.
@@
Figure 6: Comparing Standard Network Features and Ricci Curvature. For a select set of states in the Philadelphia (green), Chicago (purple), and Denver (gray) HSA regions, we visualize distributions of clustering coefficients (CC) in comparison to $\kappa_{\text{OR}}$. These are different measures (with clearly different distributions) for capturing connectivity in a network. In the context of care delivery positive CC and $\kappa_{\text{OR}}$ values together could indicate high-functioning subregions of a Physician Referral Network.
Additionally, we provide distributions of Forman-Ricci and Ollivier-Ricci Curvature values for network edges, grouped by state and region in Figures 4 and 5. As we observe from the distribution, ORC values are constrained to a much tighter range, with the most negatively curved edges in our distribution approaching a value of $-2$, while all edge curvatures are bounded above by $+1$. In comparison, Forman-Ricci, due to its dependence on node degree and the number of triangles, can produce very large curvature values, both negative and positive. Given the known trade-off between locality, expressivity, and efficiency between Forman-Ricci and Ollivier-Ricci Curvature [35], further analysis is required to understand which descriptors are most useful for specific healthcare management tasks. Despite clear agreement between various classical network science features (e.g, Figure 6), Ricci curvature measures are known to be much more expressive in capturing network structure. Moreover, there is precedence for alleviating bottlenecks in graph neural networks by rewiring negatively curved edges in a network [37]. This provides a novel way to potentially “rewire” care delivery networks using these measures in order to improve healthcare quality in regions that suffer from care fragmentation.
As part of our preliminary analysis, we compared Forman-Ricci Curvature with various population statistics, including Medicare enrollment, total mortality, and the proportion of non-white populations. As seen in Figure 7, our findings reveal that large networks exhibiting low average curvature—indicating many potential bottlenecks in care delivery—correlate with high Medicare enrollment, elevated mortality rates, and higher proportions of Black and Hispanic populations. These results (which can be reproduced interactively with apparent) suggest that negatively curved edges are indicative of structural properties in physician networks commonly found in less affluent HSAs.